Circular No. 8122

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 H1 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been reported to be cometary on Apr. 25 CCD frames taken by H. Mikuž (Crni Vrh, 0.60-m reflector + R filter; strongly condensed with coma diameter ∼ 20′′ and m_{1} = 15.9), P. Kušnirák (Ondřejov, 0.65-m reflector; “seems to be slightly diffuse”), and T. Spahr (Mount Hopkins, 1.2-m reflector; faint fan-shaped tail ∼ 5′′ long toward the south).

2003 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 24.38253  18^{h}15^{m}48.^{s}00  +19^{o}55′25′′.3  16.9

Astrometry, very preliminary parabolic orbital elements (T = 2004 Mar. 5, q = 2.86 AU, i = 140^{o}), and an ephemeris appear on MPEC 2003-H28.

COMET C/2003 H2 (LINEAR) Another apparently asteroidal object reported by LINEAR (discovery

observation below), and posted on the NEO Confirmation Page, has also been reported to be cometary on CCD frames taken on Apr. 25 by Mikuž (diffuse with condensation and coma diameter ∼ 20′′), M. Tichý (Kleť, 1.06-m reflector; diffuse with faint tail in p.a. 270^{o}), and Kušnirák (coma diameter ∼ 10′′).

2003 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 24.39543  18^{h}27^{m}01.^{s}22  +22^{o}49′33′′.4  18.8

Astrometry, very preliminary parabolic orbital elements (T = 2003 May 4, q = 2.25 AU, i = 74^{o}), and an ephemeris appear on MPEC 2003-H29.

SUPERNOVA 2003du IN UGC 9391 R. Kotak and W. P. S. Meikle, Imperial College, London, on behalf of

the European Supernova Collaboration, report that a spectrum (range ∼ 500–760 nm) of SN 2003du (cf. IAUC 8121), obtained at the Isaac Newton Telescope (+ IDS spectrograph) by P. Rodriguez-Gil on Apr. 24.06 UT, shows the characteristic Si II 615-nm absorption feature indicating that it is of type Ia. Comparison with the spectra of SN 2002bo indicates a phase of ≈ 2 weeks before maximum light (−13 ± 2 days). The depths of the features, relative to the continuum, appear somewhat more shallow than in SN 2002bo.

2003 April 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green